Joe has a collection of nickels and dimes that is worth $6.05. If the number of dimes was doubled and the number of nickels was decreased by 10, the value of the coins would be $9.85. How many dimes does he have?
Answer by nerdybill(7384)

Joe has a collection of nickels and dimes that is worth $6.05. If the number of dimes was doubled and the number of nickels was decreased by 10, the value of the coins would be $9.85. How many dimes does he have?
.
Let d = number of dimes
and n = number of nickels
.
From:"Joe has a collection of nickels and dimes that is worth $6.05."
.10d + .05n = 6.05 (equation 1)
.
From:"If the number of dimes was doubled and the number of nickels was decreased by 10, the value of the coins would be $9.85"
.10(2d) + .05(n-10) = 9.85
.20d + .05n-.5 = 9.85
.20d + .05n = 10.35 (equation 2)
.
Solve equation 1 for n:
.10d + .05n = 6.05
.05n = 6.05 - .10d
n = 121 - 2d
.
Plug the above into equation 2 and solve for d:
.20d + .05n = 10.35
.20d + .05(121 - 2d) = 10.35
.20d + 6.05 - .1d = 10.35
.10d + 6.05 = 10.35
.10d = 4.3
d = 43 dimes